TMS Meeting Minutes, December 9, 2021

Attendees: Andrew, Breann, Christian, Heejoo, John, Olga, Trenton, Yang (Zoom Meeting)

TMS Operations:
  • There were no opportunities to run TMS recently
  • Heejoo was able to test the USB cable functioning for five hours using the command window
    • When the cable was malfunctioning, there were restricted access errors in the command window
  • It will be great when this is working properly and
  • John made plots of all recent TMS operation, and Heejoo uploaded these plots on the Google Drive, and included all email correspondence
H/W Updates:
  • Fiber inspection device is held up in the worldwide shipping fiasco
  • Channels 10 and 11 still need some work
  • Heejoo and Trenton to head up soon to work on the channels and the collimator mounts
  • Heejoo preparing a fiber integrity report to quantify the health of our fibers
  • Need to re-up the discussion with Etalon about purchasing a MPO fiber box for multi-core fiber operations
S/W Updates:
  • Yang updated the TMS code to include the thermal Z4 correction in the log files
Commissioning Document:
  • Looks like we are getting towards a robust, hands-off system that is ready for operation on-sky
  • There are a few things that are not quite ready
    • Channel dropping is still a problem, and we don't have a perfect solution yet
      • We have channel redundancy, so dropping shouldn't be a problem, but we don't drop gracefully as of yet
      • Andrew suggests using the pose acceleration calculation to handle this dropping well
      • Supply a correction term to the new channel config w/ dropped channels that is constructed using a good reference
      • With an 8 channel measurement that's good, then use that to construct the correction once a channel is dropped as a constant correction applied until the channel is regained
      • Last time we did this, we had some issues with dropping, so we have room to improve on this
        • When a new measurement comes with a missing channel, the software goes back and looks for the last good measurement
        • The software computes a pose with the last good measurement, and a difference is taken between the pose calculated with the last good measurement and the most recent measurement with a dropped channel
        • This difference is continuously subtracted from the new, reduced channel configuration
      • ~4 micron uncorrelated channel noise appears to be causing much of the problem
      • Andrew requests that we have the ability to toggle channel dropping mode on and off for debugging
    • The thermal correction needs improvement, as well
      • Andrew suggests using the GMT report on the thermal analysis
      • Unfortunately, there are temperature sensors that are missing, so we will need to consider how to best handle the non-uniform distribution (consider using weighting and gain factors)
      • When there are particularly large temperature changes on the mirror, the thermal correction doesn't work well
      • Andrew is convinced that there is a borosilicate mirror tilt effect that is ruining the pointing
      • We also really need to dig dipper into the thermal focus correction
  • We are looking to get another engineering night as per Christian's suggestion to validate TMS
    • Night of the 26th and 28th of December, we will have engineering time
    • The priority is for other tasks, but we should have enough time to move to TMS after this
    • Christian to work with Olga to prepare a sequence of observations to test TMS, active vs. passive per side
    • We may have just a few hours, but if we have a moving target, we can get some good information
    • Beyond this, it could be tough to get something before the end of January, but Christian to start preparing the schedule for next semester
    • There is a pool of engineering and discretionary time
    • We may have some issues with trying to run active TMS on one side, and passive TMS on the other
    • We don't really have a good way to do this, so we will work out a better strategy
  • Andrew proposes offering a shared risk mode of TMS operation to observers, noting that TMS does improve the outcome of LBC operation
    • Christian agrees, partially, but wants to look at the operational side, what is still needed, what information must be accessible
    • Do we need someone on-call from the TMS team, or do we have what we need right now?
    • Probably need a bit more clarity in the wiki and the commissioning documents, but we are pretty darn close right now
    • Andrew notes that we also really need the TMS co-pointing model operational on LBC
    • John raises an issue with this, however; the current non-TMS pointing model is designed to work with mixed instruments while co-pointing (i.e. LBCr + MODS), though this is relatively rare; this means you can't co-point the two sides, as they differ by a few arcmin
    • For now, we will need to hold on to the old pointing model, and develop the new one soon
SPIE Deadline January 5th:
  • Andrew to prepare a draft of an abstract

-- TrentonBrendel - 09 Dec 2021
Topic revision: r1 - 09 Dec 2021, TrentonBrendel
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback